Find the nature of the planar section of the conicoid x^2/3 −y^2/4 = z by the plane x+2y−z = 6
1
Expert's answer
2020-05-18T00:57:30-0400

"\dfrac{x^2}{3}-\dfrac{y^2}{4}=z"


"x+2y\u2212z=6"

We will find z

"z=x+2y\u22126"


"4\n\nx^\n2\n\u200b\n\n\u2212\n3\n\ny^\n2\n\u200b\n\n=12(x+2y\u22126)"

"4(x^\n2\n\u22123x)\u22123(y^\n2\n+8y)=\u221272"


"4(x−\frac{3}{2})
^2
−3(y+4)
^2
=−72+9−48=−111"


"4\n\n\\frac{\n\u200b\n\n\n(x\u2212\\frac\n\n\n32\n\u200b\n\n)^2}{111}\n\u200b\n\n\u2212\n3\n\\frac{\n\n\u200b\n\n\n(y+4)^\n2}{111}\n\u200b\n\n=\u22121"


"\\frac{\n\u200b\n\n\n(x\u2212\\frac\n\n\n32\n\u200b\n\n)^2}{\\frac{111}{4}}\n\u200b\n\n\u2212\n\n\\frac{\n\n\u200b\n\n\n(y+4)^\n2}{\\frac{111}{3}}\n\u200b\n\n=\u22121"


This is a equation of conjugated hyperbola

"center(\n\\frac\n\n32\n\u200b\n\n,\u22124)"

semiaxes=

"a=\n\\sqrt{\n\n111\/4\n\u200b}\n\n\n\u200b\n\n,b=\n\\sqrt{\n\n111\/3\n\u200b\n\n}\n\u200b"